A pack of 52 cards were distributed equally among 4 players. Find the chance that 4 kings are
held by particular player.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
When a pack of 52 cards were distributed equally among 4 players,the chance that 4 kings are held by particular player is given by
=13/52 x 12/51 x 11/50 x 10/49 = 11/4165
Thanks and Regards
Shaik Aasif
askIITians faculty